### v2

# Unfitted finite element interpolated neural networks

Abstract.
We present a novel approach that integrates unfitted finite element methods and neural networks to approximate partial differential equations on complex geometries. Easy-to-generate background meshes (e.g., a simple Cartesian mesh) that cut the domain boundary (i.e., they do not conform to it) are used to build suitable trial and test finite element spaces. The method seeks a neural network that, when interpolated onto the trial space, minimises a discrete norm of the weak residual functional on the test space associated to the equation. As with unfitted finite elements, essential boundary conditions are weakly imposed by Nitsche’s method. The method is robust to variations in Nitsche coefficient values, and to small cut cells. We experimentally demonstrate the method’s effectiveness in solving both forward and inverse problems across various 2D and 3D complex geometries, including those defined by implicit level-set functions and explicit stereolithography meshes. For forward problems with smooth analytical solutions, the trained neural networks achieve several orders of magnitude smaller $H^{1}$ errors compared to their interpolation counterparts. These interpolations also maintain expected $h$- and $p$-convergence rates. Using the same amount of training points, the method is faster than standard PINNs (on both GPU and CPU architectures) while achieving similar or superior accuracy. Moreover, using a discrete dual norm of the residual (achieved by cut cell stabilisation) remarkably accelerates neural network training and further enhances robustness to the choice of Nitsche coefficient values. The experiments also show the method’s high accuracy and reliability in solving inverse problems, even with incomplete observations.
Key words and phrases:
neural networks, PINNs, unfitted finite elements, PDE approximation, complex geometries, inverse problems
1. Introduction
[…]
pde are powerful mathematical tools for modelling a wide range of physical phenomena, including heat transfer, fluid dynamics, and electromagnetics. Since obtaining closed-form analytic solutions to these equations is impractical (if not unfeasible) in most cases, numerical methods like the FEM are often employed to compute approximate solutions instead. This method makes use of piecewise polynomial spaces on a partition of the computational domain to approximate the solution of the partial differential equation (PDE). Its many advantages, such as a solid mathematical foundation [Ern2021], the capability to handle complex geometries, and the ease of imposing boundary conditions, render FEM a widely adopted approach for computationally solving PDEs.
Problems posed on complex and/or evolving geometries frequently arise in a wide range of practical applications, such as those modelled by free-boundary problems. The use of unstructured mesh generators to generate body-fitted meshes that conform to such geometries is often impractical, challenging, time-consuming, and requiring human intervention in many cases. Unfitted FEMs, also known as embedded or immersed boundary FEMs, are especially compelling in these scenarios. The differential equation at hand is discretised by embedding the computational domain in an easy-to-generate background mesh (e.g., a simple Cartesian mesh) that does not necessarily conform to its geometrical boundary, thus drastically reducing the geometrical constraints imposed on the meshes to be used for discretisation. However, unfitted FEMs introduce their own numerical challenges, such as the small cut cell problem, where arbitrarily small intersections between the background mesh and the physical domain can lead to numerical instability (i.e., arbitrarily ill-conditioning of the system resulting from discretisation). We refer the reader to [dePrenter2023] for a comprehensive and updated review of unfitted FEMs. Two prominent unfitted FEMs, that we leverage in this work, CutFEM [Burman2015] and aggregated (AgFEM[Badia2018], address this issue in different ways: CutFEM introduces a ghost penalty term [Burman2010] to stabilise the system, while AgFEM eliminates ill-posed cells through aggregation techniques and discrete extension operators. CutFEM and AgFEM have been successfully applied to a wide range of PDEs on complex geometries, e.g., interface problems [Neiva2021], fluid-structure interaction [Massing2015], and linear elasticity [Hansbo2017].
[…]
Several studies in the literature explore the use of NNs to solve PDEs on complex geometries. PhyGeoNet [Gao2021] addresses this challenge by mapping the irregular physical domain to a regular reference domain, where classic convolutional s can be trained to approximate the PDE solution. However, PhyGeoNet lacks the flexibility to handle highly complex domains, as establishing a one-to-one mapping for such geometries is difficult, and direct application of elliptic coordinate transformation is often impossible. Another popular approach involves training one NN in the interior domain and prescribing an offset function that either satisfies or is trained to satisfy the Dirichlet boundary conditions. The final solution is obtained by adding the interior NN, multiplied by a distance function that vanishes on the Dirichlet boundary, to the offset function [Berg2018, Sheng2021]. The shortcoming of this method is that neither the offset function nor the distance function is easily obtainable for complex geometries. The $\Delta$-PINN method [SahliCostabal2024] enables the use of PINNs on complex geometries by representing manifolds with eigenfunctions of the Laplace-Beltrami operator. These functions can be approximated numerically for any shape using FEs. When applied to a steady-state heat transfer problem on a complex sink, the method produces a temperature distribution that is qualitatively closer to the ground truth compared to standard PINNs. However, the number of eigenfunctions has a significant, non-monotonic impact on the accuracy of the method, and must be carefully tuned for complex geometries.
[…]
The unfitted FEINN method transforms the (non)linear system of equations into a residual minimisation problem. The method is insensitive to small cut cell issues when residual vector-based loss functions (that is, based on algebraic norms applied to the residual vector) are employed. Even with loss functions based on the dual norm of the residual functional, only the computation of the discrete dual norm of the residual, i.e., an approximation of the $H^{1}$ inner product for the linearised FE space, requires stabilisation; this is a simpler task compared to stabilising the original unfitted problem with a high-order trial space and a linearised test space. A comprehensive set of experiments on 2D and 3D complex geometries demonstrates the effectiveness of unfitted FEINNs. We evaluate the method on linear and nonlinear forward PDEs, as well as inverse nonlinear problems, to confirm its reliability for solving PDEs with weak $H^{1}$ stability. The geometries considered in the experiments include implicitly defined ones via level-set functions, as well as those defined explicitly by stereolithography (STL) meshes, which are more representative of real-world applications [Badia2022b]. The method is compared against standard PINNs on both CPU and GPU architectures. Employing the same training points, unfitted FEINNs reach comparable accuracy using fewer training iterations and reduced computational time. The trained NNs significantly outperform FE interpolations of the analytic solution in terms of $H^{1}$ error, particularly when the solution is smooth. Furthermore, when a discrete dual norm of the residual is applied, we achieve further gains in accuracy, and the number of iterations required for convergence is substantially reduced.
The rest of the article is organised as follows. Sect. 2 introduces the continuous setting, unfitted FEs, the unfitted FEINN method, and their associated loss functions. Sect. 3 details the implementation of linearised test FE spaces and provides an overview of the FEINN implementation. Sect. 4 presents numerical experiments with different PDEs on 2D and 3D complex geometries. Finally, Sect. 5 concludes the article and outlines promising directions for future research.
2. Methodology
2.1. Continuous setting
Let $\Omega\subset\mathbb{R}^{d}$ be a Lipschitz polyhedral domain with boundary $\partial\Omega$ and outward pointing unit normal $\boldsymbol{n}$. We focus on low-dimensional problems, i.e., $d\in\{2,3\}$. The physical domain $\Omega$ may have a complex shape represented implicitly by a level set function $\phi:\mathbb{R}^{d}\to\mathbb{R}$ such that $\Omega=\{\boldsymbol{x}\in\mathbb{R}^{d}:\phi(\boldsymbol{x})<0\}$ and $\partial\Omega=\{\boldsymbol{x}\in\mathbb{R}^{d}:\phi(\boldsymbol{x})=0\}$. Alternatively, the boundary $\partial\Omega$ can be explicitly described using STL surface meshes [Badia2022b].
As a model problem, we consider the Poisson equation with Dirichlet boundary conditions. Its strong form reads: find $u\in U\doteq H^{1}(\Omega)$ such that
$-\boldsymbol{\Delta}u=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$ (1)
where $f\in H^{-1}(\Omega)$ is a known source term, and $g\in H^{1/2}(\partial\Omega)$ is a known Dirichlet boundary data. Note that Neumann boundary conditions can also be included into the formulation; however, for simplicity and without loss of generality, we focus on the pure Dirichlet case. The Poisson equation is a representative example of second-order elliptic PDE with weak $H^{1}$ stability, which is the class of PDEs considered in this work. Besides, we adopt a nonlinear PDE from [Berrone2022] that fits in this class. The strong form of the nonlinear PDE is as follows: find $u\in U$ such that
$-\boldsymbol{\Delta}u+(\boldsymbol{\beta}\cdot\boldsymbol{\nabla})u+\sigma\mathrm{e}^{-u^{2}}=f\quad\text{in }\Omega,\qquad u=g\quad\text{on }\partial\Omega,$ (2)
where $\boldsymbol{\beta}$ and $\sigma$ are known coefficients. We emphasise that the methodology in this work can be readily extended to more general PDEs, as far as suitable unfitted FE discretisation schemes are available.
2.2. Unfitted finite element discretisation
Since $\Omega$ has a complex shape and generating a mesh that fits $\Omega$ is typically challenging, we use unfitted FEs to discretise the space. We introduce a background box $\Omega^{\rm box}$ that contains $\Omega$ (i.e., $\Omega\subset\Omega^{\rm box}$), and can be easily partitioned by mean of an easy-to-generate mesh, e.g., a simple Cartesian mesh. Consider a quasi-uniform triangulation $\mathcal{T}_{h}$ of $\Omega^{\rm box}$ with the characteristic mesh size $h>0$. Following the notation in [Badia2018], a cell $K\in\mathcal{T}_{h}$ can belong to one of three categories: if $K\cap\Omega=\emptyset$, then $K$ is an external cell and $K\in\mathcal{T}_{h}^{\rm ext}$; if $K\cap\Omega=K$, then $K$ is an interior cell and $K\in\mathcal{T}_{h}^{\rm in}$; otherwise, $K$ is a cut cell and $K\in\mathcal{T}_{h}^{\rm cut}$. The set of active cells is defined as $\mathcal{T}_{h}^{\rm act}\doteq\mathcal{T}_{h}^{\rm in}\cup\mathcal{T}_{h}^{\rm cut}$. We then define the standard FE space $U_{h}^{\rm std}\subset U$ of order $k_{U}$ on $\mathcal{T}_{h}^{\rm act}$. Note that $U_{h}^{\rm std}$, a conforming Lagrangian FE space on interior and cut cells, is exposed to ill-conditioning due to small cut cells [Burman2015, Badia2018].
Unfitted FEs mainly face three challenges: (i) numerical integration over the physical domain $\Omega$; (ii) the imposition of essential boundary conditions; and (iii) stability issues caused by small cut cells. We give details on how to address (i) in Sect. 3.1. To address (ii), Nitsche’s method is widely used to weakly enforce Dirichlet boundary conditions, and we use it in this work as well to build NNs such that their interpolation onto the trial FE space weakly satisfy the essential boundary conditions. In the following sections, we first introduce Nitsche’s method and then discuss two unfitted FE formulations to address (iii).
2.3. Nitsche’s method
With unfitted space discretisations, strongly enforcing essential boundary conditions in the approximation space (as it is done with body-fitted FEs) is challenging. Instead, we use Nitsche’s method to weakly impose these conditions. We denote the trial and test FE spaces as $U_{h}$ and $V_{h}$, respectively. Note that $U_{h}\subseteq U$ and $V_{h}\subseteq U$ can either be the standard FE spaces introduced in Sect. 2.2 or the aggregated FE spaces introduced in Sect. 2.4. When necessary, the type of space is indicated with a superscript. We now define the bilinear form $a:U_{h}\times V_{h}\to\mathbb{R}$ and the linear form $\ell:V_{h}\to\mathbb{R}$ for the Poisson model problem (1) as
$a(u_{h},v_{h})$ $\doteq\int_{\Omega}\boldsymbol{\nabla}u_{h}\cdot\boldsymbol{\nabla}v_{h}\mathrm{d}V+\int_{\partial\Omega}(\frac{\gamma}{h}u_{h}v_{h}-u_{h}\frac{\partial v_{h}}{\partial\boldsymbol{n}}-\frac{\partial u_{h}}{\partial\boldsymbol{n}}v_{h})\mathrm{d}S,$
$l(v_{h})$ $\doteq\int_{\Omega}fv_{h}\mathrm{d}V+\int_{\partial\Omega}(\frac{\gamma}{h}gv_{h}-g\frac{\partial v_{h}}{\partial\boldsymbol{n}})\mathrm{d}S,$
where $\gamma>0$ is the Nitsche coefficient and $\partial/\partial\boldsymbol{n}$ denotes derivative in the direction of $\boldsymbol{n}$. These forms are derived by multiplying the Poisson equation (1) by a test function $v_{h}\in V_{h}$, integrating by parts while noting that $v_{h}\neq 0$ on $\partial\Omega$, and adding the Nitsche boundary terms. Then, the Nitsche FE formulation reads: find $u_{h}\in U_{h}$ such that
$a(u_{h},v_{h})=\ell(v_{h}),\quad\forall v_{h}\in V_{h}.$ (3)
In FEM, the Nitsche coefficient $\gamma$ in $a$ and $\ell$ is a mesh-dependent parameter that must be sufficiently large to guarantee the coercivity of the bilinear form. Two suggestions to determine $\gamma$ are summarised in [Burman2015, Sect. 2.1]. It is straightforward to verify that Nitsche’s method is consistent with the original problem: when $u_{h}=g$ on $\partial\Omega$, the Nitsche boundary terms in $a$ and $\ell$ vanish, and the formulation reduces to the integration-by-parts form of (1). In regards to the convergence of Nitsche’s method, we refer to [Thomee2007, Chap. 2].
We denote the weak FE residual functional as
$\mathcal{R}_{h}(u_{h})\doteq a(u_{h},\cdot)-\ell(\cdot)\in V_{h}^{\prime}.$ (4)
2.4. CutFEM and AgFEM
Both CutFEM and AgFEM are unfitted methods that use the FE formulation (3), the difference lies in the way they stabilise the FE scheme. CutFEM [Burman2015] utilises the standard FE spaces $U_{h}^{\rm std}$ and $V_{h}^{\rm std}$ introduced in Sect. 2.2. To address the small cut cell problem, the CutFEM adds ghost penalty terms [Burman2010] to the left-hand side of (3). The penalty is defined on the skeleton $\mathcal{F}_{g}$, which is the set of interior faces that belong to an element $K$ such that $K\cap\partial\Omega\neq\emptyset$ [Hansbo2017]. Let us denote the normal to a face $F\in\mathcal{F}_{g}$ as $\boldsymbol{n}_{F}$, the ghost penalty terms are defined as
[…]
Since high-order directional derivative terms are required in the ghost penalty (5), CutFEM becomes computationally expensive when $k_{U}$ is high.
Instead of adding additional terms to the formulation, AgFEM [Badia2018] overcomes this issue at the FE space level by cell aggregation and a discrete extension operator. Specifically, ill-posed cut cells are absorbed by well-posed interior cells to form aggregates, with each ill-posed cut cell belonging to only one aggregate, and each aggregate containing only one well-posed interior cell, i.e., the so-called root cell of the aggregate. The aggregated FE spaces $U_{h}^{\rm ag}$ and $V_{h}^{\rm ag}$ are then defined on the aggregated mesh $\mathcal{T}_{h}^{\rm ag}$. In such FE spaces, the degrees of freedom in each aggregate are associated only with the root cell, while the other ill-posed DoFs are linear combinations of those in the root cell (the discrete extension from interior to cut cells). We refer to, e.g., [Badia2018], for details of the cell aggregation algorithms, analysis, and implementation details of this method.
2.5. Neural networks
We focus on fully-connected, feed-forward NNs composed of affine maps followed by nonlinear activation functions. The network structure is represented by a tuple $(n_{0},\ldots n_{L})\in\mathbb{N}^{L+1}$, where $L$ is the number of layers and $n_{k}$ is the number of neurons at layer $1\leq k\leq L$. We set $n_{0}=d$ for the input layer and $n_{L}=1$ for the output layer, and $n$ neurons for all hidden layers, i.e., $n_{1}=n_{2}=...=n_{L-1}=n$.
At each layer $1\leq k\leq L$, the affine map is represented by $\boldsymbol{\Theta}_{k}:\mathbb{R}^{n_{k-1}}\to\mathbb{R}^{n_{k}}$ and defined as $\boldsymbol{\Theta}_{k}\boldsymbol{x}=\boldsymbol{W}_{k}\boldsymbol{x}+\boldsymbol{b}_{k}$, where $\boldsymbol{W}_{k}\in\mathbb{R}^{n_{k}\times n_{k-1}}$ is the weight matrix and $\boldsymbol{b}_{k}\in\mathbb{R}^{n_{k}}$ is the bias vector. After the application of the affine map corresponding to each layer except the last one, the activation function $\rho:\mathbb{R}\to\mathbb{R}$ is applied element-wise. With these definitions, the network is a parametrisable function $\mathcal{N}(\boldsymbol{\theta}):\mathbb{R}^{d}\to\mathbb{R}$ defined as:
$\mathcal{N}(\boldsymbol{\theta})=\boldsymbol{\Theta}_{L}\circ\rho\circ\boldsymbol{\Theta}_{L-1}\circ\ldots\circ\rho\circ\boldsymbol{\Theta}_{1},$
where $\boldsymbol{\theta}$ represents all the trainable parameters $\boldsymbol{W}_{k}$ and $\boldsymbol{b}_{k}$. We apply the same fixed activation function at each layer, even though the activations could be different or trainable. In this work, $\mathcal{N}$ denotes the NN architecture, and $\mathcal{N}(\boldsymbol{\theta})$ represents a particular realisation of this network.
2.6. Unfitted finite element interpolated neural networks
The unfitted FEINN method extends the FEINN formulation in [Badia2024] to solve elliptic PDEs on complex domains by using unfitted FE formulations (3). We define a FE interpolant $\pi_{h}:\mathcal{C}^{0}\rightarrow U_{h}^{\rm std}$. For the grad-conforming Lagrange (nodal) spaces considered in this work, $\pi_{h}$ involves only pointwise evaluations at the free nodes. The unfitted FEINN method aims to find:
$u_{\mathcal{N}}\in\underset{w_{\mathcal{N}}\in\mathcal{N}}{\mathrm{arg\,min}}\,\mathscr{L}(w_{\mathcal{N}}),\qquad\mathscr{L}(w_{\mathcal{N}})\doteq\left\lVert\mathcal{R}_{h}({\pi}_{h}(w_{\mathcal{N}}))\right\rVert_{X^{\prime}},$ (6)
where $u_{\mathcal{N}}$ is the NN approximation to the solution of the problem, and $\left\lVert\cdot\right\rVert_{X^{\prime}}$ denotes an appropriate discrete dual norm.
Let us now discuss the choice of the trial and test FE spaces. Since the FE problem (3) is reformulated as a minimisation problem (6), and the trial FE space is “hidden” in the PDE residual (4), we can choose the trial space to be $U_{h}^{\rm std}$, i.e., the standard unfitted FE space without any stabilisation. It is not necessary to apply stabilisation because the discrete minimisation problem is determined by the NN parameters, not the ill-posed DoFs in the FE space. In fact, all experiments in Sect. 4 utilise $U_{h}^{\rm std}$ as the trial space and yield accurate results. When NNs are interpolated onto $U_{h}^{\rm ag}$, they lose control over the cut cells during training because these DoFs are defined as linear combinations of the DoFs in the root cells. For the test FE space in the residual $\mathcal{R}_{h}$ (4), the choice depends on the specific loss function employed. Options include $V_{h}^{\rm std}$, with or without a ghost penalty, or $V_{h}^{\rm ag}$. Details on the loss functions and their corresponding test space selections are provided in Sect. 2.7. Moreover, as shown in [Badia2024, Badia2025adaptive], the linearisation of test spaces significantly accelerates convergence and improves the accuracy of NN solutions. In unfitted FEINNs, we also employ such Petrov-Galerkin FE formulation, adopting linearised $V_{h}^{\rm std}$ and $V_{h}^{\rm ag}$. These spaces are constructed by refining the background mesh discretising $\Omega^{\rm box}$ and building linear FE spaces on the active mesh resulting from intersecting the refined mesh and the domain boundary. Implementation details of these linearised unfitted test spaces are discussed in Sect. 3.1.
2.7. Loss functions
The FE residual $\mathcal{R}_{h}$ (4) is isomorphic to the residual vector
$[\mathbf{r}_{h}(w_{h})]_{i}=\left<\mathcal{R}_{h}(w_{h}),\varphi^{i}\right>\doteq\mathcal{R}_{h}(w_{h})(\varphi^{i}),$
where $\{\varphi^{i}\}_{i=1}^{N}$ are the shape functions that span the test FE space $V_{h}$ and $N$ is the dimension of $V_{h}$. We can use the following loss function:
$\mathscr{L}(u_{\mathcal{N}})=\left\lVert\mathbf{r}_{h}(\pi_{h}(u_{\mathcal{N}}))\right\rVert_{\chi},$ (7)
where $\left\lVert\cdot\right\rVert_{\chi}$ is an algebraic norm, e.g., $\ell^{1}$ or $\ell^{2}$ norm. When using a body-fitted mesh and the $\ell^{2}$ norm, the method becomes equivalent to the Nitsche variant of IVPINNs [Berrone2023]. However, in contrast to [Berrone2023], in this work we focus on unfitted meshes to handle complex geometries. When the loss (7) is employed for unfitted FEINNs, the method is referred to as $\ell^{1}$- or $\ell^{2}$-FEINNs, depending on the algebraic norm chosen. Although the $\ell^{2}$ norm is more common in the literature [Berrone2022], the $\ell^{1}$ norm is also effective and can produce remarkable results, as revealed in [Badia2024, Badia2025adaptive]. We indeed use both norms in the experiments of this work. It is important to note that we always choose $V_{h}^{\rm std}$ as the test space when loss (7) is used. In this setup, neither the trial nor test spaces are stabilised, nor is there a need for ghost penalty terms to stabilise the discrete variational FE formulation. Nevertheless, as demonstrated in the moving domain tests in Sect. 4.1.3, the unfitted FEINN method remains robust to small cut cells. In this setting, the problem is stated as a residual minimisation, where several (local) minima do not disrupt the algorithm. This contrasts sharply with unfitted FEM, where the solution requires multiplying the inverse of the system matrix and the right-hand side vector, facing ill-conditioning issues, i.e., a lack of uniqueness in the singular limit.
However, as mentioned in [Badia2024, Badia2025adaptive], the loss (7) becomes ill-posed as $h\rightarrow 0$; it amounts to computing the $L^{2}$-norm (up to a constant that depends on the quasi-uniformity of the mesh) of the residual functional, which is only in $H^{-1}(\Omega)$. Besides, using these algebraic norms on the residual vector, optimal convergence rates have not been proven [Berrone2022]. It has motivated the use of dual norms in [Badia2024], i.e.,
$\mathscr{L}(u_{\mathcal{N}})=\left\lVert\mathcal{R}_{h}(\pi_{h}(u_{\mathcal{N}}))\right\rVert_{U^{\prime}}.$ (8)
In FEINNs, this discrete dual norm can be computed by introducing a discrete Riesz projector $\mathcal{B}_{h}^{-1}:V_{h}^{\prime}\to V_{h}$ such that
$\mathcal{B}_{h}^{-1}\mathcal{R}_{h}(w_{h})\in V_{h}\ :\left(\mathcal{B}_{h}^{-1}\mathcal{R}_{h}(w_{h}),v_{h}\right)_{U}=\mathcal{R}_{h}(w_{h})(v_{h}),\quad\forall v_{h}\in V_{h}.$
where $\mathcal{B}_{h}$ is (an approximation of) the corresponding Gram matrix in $V_{h}$. Then, we can write the loss preconditioned by $\mathcal{B}_{h}$ as:
$\mathscr{L}(u_{\mathcal{N}})=\left\lVert\mathcal{B}_{h}^{-1}\mathcal{R}_{h}({\pi}_{h}(u_{\mathcal{N}}))\right\rVert_{X}.$ (9)
When $X=U$, the loss functions in (8) and (9) are equivalent, and the method is denoted as $H^{-1}$-FEINNs. Other choices of $X$ are possible, e.g., $X=L^{2}(\Omega)$. To reduce computational costs, $\mathcal{B}_{h}$ can be substituted with any spectrally equivalent approximation; for example, geometric multigrid (GMG) preconditioners are explored in [Badia2024].
When utilising the loss function with the discrete dual norm of the residual (9), the test FE space can be either $V_{h}^{\rm std}$ or $V_{h}^{\rm ag}$. However, adding the ghost penalty (5) is essential to stabilise the variational problem (9) in $V_{h}^{\rm std}$, ensuring its invertibility in inexact arithmetic regardless of the small cuts. In contrast, discretising (9) in $V_{h}^{\rm ag}$ is inherently stable [Badia2018]. Indeed, in [Verdugo19], it was shown that standard multigrid solvers for body-fitted FEM reach expected mesh-independent convergence rates and linear complexity when applied to the Gram matrix in $V_{h}^{\rm ag}$. We compare the performance of these two test spaces in the experiments in Sect. 4.1.5.
The analysis of FEINNs for the loss function in (8) can be found in [Badia2025adaptive, Sect. 3] for strong imposition of boundary conditions and coercive systems. The extension to Nitsche’s method is straightforward, with minor modifications. These results can be used in the case of unfitted FEINNs under the assumption of a discrete inf-sup condition between the trial and test spaces (see [Badia2025adaptive, Sect. 2.2]) combined with the enhanced stability properties of CutFEM and AgFEM discretisation [Burman2010, Badia2018].
2.8. Extension to inverse problems
Unfitted FEINNs can be extended straightforwardly to solve inverse problems. These include estimating the unknown model parameters of a PDE from partial and/or noisy observations of the solution and fully reconstructing the solution/state itself. Let us denote the unknown parameters as $\boldsymbol{\Lambda}$, which may include physical coefficients, boundary values, or source terms. In addition to the state NN $u_{\mathcal{N}}$, we introduce another NNs $\boldsymbol{\Lambda}_{\mathcal{N}}$ to approximate the unknown model parameters. We interpolate $\boldsymbol{\Lambda}_{\mathcal{N}}$ onto suitable FE spaces using the interpolants $\boldsymbol{\pi}_{h}$ to ensure efficient computation of the loss function and its gradients. The loss function for the inverse problem is defined as:
$\mathscr{L}(\boldsymbol{\Lambda}_{\mathcal{N}},u_{\mathcal{N}})\doteq\left\lVert\mathbf{d}-\mathcal{D}(u_{\mathcal{N}})\right\rVert_{\ell^{2}}+\alpha\left\lVert\mathcal{R}_{h}(\boldsymbol{\pi}_{h}(\boldsymbol{\Lambda}_{\mathcal{N}}),\pi_{h}(u_{\mathcal{N}}))\right\rVert_{X^{\prime}},$ (10)
where $\mathbf{d}\in\mathbb{R}^{M}$ is the observed data and $M\in\mathbb{N}$ is the number of observations, $\mathcal{D}:U\rightarrow\mathbb{R}^{M}$ is the observation operator, and $\alpha\in\mathbb{R}^{+}$ is the penalty coefficient. The first term in (10) measures the data misfit, while the second term weakly imposes the PDE constraint. In the inverse problem experiments in Sect. 4.3, $\mathcal{D}$ is defined by directly evaluating $u_{\mathcal{N}}$ at the observation points. We refer to [Badia2024] for a detailed discussion on FEINNs for inverse problems. The key difference in this work is the use of unfitted FE formulations rather than body-fitted FE formulations.
3. Implementation
Unfitted FEINNs are a natural extension of the FEINNs framework, thus allowing one to reuse most of the building blocks of the software implementing the framework. The additional components required by unfitted FEINNs are: (i) the design and implementation of unfitted linearised test FE spaces and (ii) the integration over the physical domain with irregular shapes when such domain is cut by a background mesh. We first address (i) and (ii) in Sect. 3.1, followed by an outline of the Julia implementation of unfitted FEINNs in Sect. 3.2.
3.1. Unfitted linearised test spaces
The (body-fitted) FEINN method in [Badia2024] constructs a trial FE space of order $k_{U}$ on a coarse mesh. It applies $k_{U}-1$ levels of uniform refinement to this coarse mesh and builds a test FE space made of linear polynomials on the refined mesh. We call such space a linearised test FE space. It has the same dimension (i.e., number of DoFs) as the trial FE space. In practice, such an approach is proven to drastically accelerate training convergence and improve NN accuracy [Badia2024].
In the unfitted case, we can use (and we indeed use) an analogous approach to construct a linearised test space. In particular, we compute a refined background mesh, and then we intersect this mesh and the physical domain. However, as shown below with an example, this can lead to a refined active triangulation that covers only a strict subregion of the coarse active triangulation. As a consequence, a test space made of linear polynomials on this refined active triangulation may result in a reduced number of DoFs compared to those in the trial space. Fortunately, the residual minimisation approach underlying unfitted FEINNs is flexible enough to naturally accommodate such scenarios.
Consider a physical domain $\Omega$ within a background box $\Omega^{\rm box}$ as displayed in Fig. 1(a). Let us assume that we discretise $\Omega^{\rm box}$ using a $5\times 4$ Cartesian mesh, which we refer to as the coarse mesh. As shown in Fig. 1(b), in this particular case, all cells in the coarse mesh are active cells. Assuming $k_{U}=2$ for the trial space $U_{h}^{\rm std}$, the global DoFs of $U_{h}^{\rm std}$ are illustrated by black dots in Fig. 1(b). Next, we uniformly refine the coarse mesh once, generating a refined mesh with $10\times 8$ cells, as shown in Fig. 1(c). The active cells on this refined mesh, resulting from the intersection of the refined mesh and the domain, are highlighted in orange. A linear test FE space $V_{h}^{\rm std}$ is then constructed on this refined active cells, with its DoFs represented by blue crosses in Fig. 1(c).
background box coarse active cells trial FE space DoFs
physical domain refined active cells test FE space DoFs
((a)) $\Omega^{\rm box}$ and $\Omega$
((b)) $U_{h}^{\rm std}$
((c)) $V_{h}^{\rm std}$
Figure 1. (a) Background box and physical domain, (b) trial and (c) test FE spaces.
We immediately observe that $V_{h}^{\rm std}$ has fewer DoFs than $U_{h}^{\rm std}$ (the same applies to $V_{h}^{\rm ag}$). A Petrov-Galerkin FE discretisation of the problem at hand results in a rectangular discrete system. While this mismatch is problematic for FEM, as it prevents the application of solvers for algebraic systems of equations, we can still apply the unfitted FEINN method because such method recasts the problem into a residual minimisation problem, which can be solved using a gradient-based optimiser. Although residual minimisation is also available in FEM (e.g., [Guermond2004]), the (body-fitted) FEINN method can produce more accurate NN solutions than FEM when the solution is smooth, as shown in [Badia2024, Badia2025adaptive, Badia2024compatible]. The computation of the discrete dual norm of the residual involves the inversion of a Galerkin discretisation operator built using the test space.
Remark 3.1.
It is actually possible to build a test FE space made of linear polynomials with the same number of DoFs as a higher order trial FE space. This requires a suitable FE on the coarse mesh active cells. In particular, one can use hierarchical Q1-iso-Qk FEs [Ern2021b, Sect. 54.4.2]. These FEs lead to a well-posed square Petrov-Galerkin discrete system as the basis functions of this FE have support across the whole coarse cell. We indeed tested this FE and observed results comparable to those achieved by linear bases on the refined active mesh resulting from cutting the refined background mesh with the domain boundary (i.e., the approach that we finally pursue in this paper). As linear bases on the refined mesh are significantly easier to implement (i.e., one can reuse verbatim existing unfitted FE codes), we chose them in this paper.
In the implementation, we first intersect the refined mesh with $\Omega$ to identify the active refined cells, which allows us to construct $V_{h}^{\rm std}$. Next, we determine the active cells on the coarse mesh using the following rule: a coarse cell is active if it has at least one active child cell. This way we do not need to compute the intersection of the coarse mesh with $\Omega$. The high-order $U_{h}^{\rm std}$ is then built on these active coarse cells. The aggregated test FE space $V_{h}^{\rm ag}$, as usual, is constructed from $V_{h}^{\rm std}$. We classify interior cells as well-posed and cut cells as ill-posed, then form aggregates with one well-posed cell as root. The space $V_{h}^{\rm ag}$ is built on these aggregates, with free DoFs determined by the root cells.
Similar to the body-fitted FEINN method, integration is performed on the refined mesh when one uses linearised FE spaces. For interior cells, the integration process is the same as in fitted FEM. For a cut cell $K$, we discretise the intersection $K\cap\Omega$ into a local-to-the-cell submesh of triangles (2D) or tetrahedra (3D) and use quadratures rule for simplices with sufficient degree to integrate exactly over these subcells those polynomials defined on the quadrilateral/hexaedral cells. This is the standard approach pursued in unfitted FEM to integrate polynomials on cut cells.
3.2. Unfitted finite element interpolated neural networks
We reuse the FEINN implementation as detailed in [Badia2024, Sect. 5] for the Euclidean norm of the residual as loss function and inverse problem cases, and in [Badia2025adaptive, Sect. 4.3] when using a discrete dual norm of the residual as loss function. For completeness, we highlight the main Julia packages used in the implementation of unfitted FEINNs. The FEM computations are managed by Gridap.jl [Badia2020, Verdugo2022], with unfitted extensions handled by GridapEmbedded.jl [GridapEmbeddedRepo] and STLCutters.jl [Martorell2024] packages. While GridapEmbedded.jl offers mesh intersection algorithms for implicit geometry representations, such as level set functions, STLCutters.jl handles explicit geometry representations using STL meshes. The neural network is implemented using Flux.jl [Innes2018]. We define custom reverse-mode rules in ChainRules.jl [ChainRules2024] for gradient computation.
4. Numerical experiments
We evaluate the accuracy of unfitted FEINNs in solving both linear and nonlinear equations on 2D and 3D complex geometries. The experiments begin with 2D and 3D forward problems, followed by inverse problems in 2D. The linear PDE considered is the Poisson equation with a constant diffusion coefficient, as described in (1). Its Nitsche FE formulation is given in (3). Note that the FEINN framework can readily handle spatially varying coefficients, as demonstrated in [Badia2024]. In the nonlinear forward and inverse experiments, we focus on the nonlinear equation in strong form (2). We first discuss the settings for forward problems.
We employ the same NN structure as in [Badia2024] for all the forward experiments, i.e., $L=5$ layers, $n=50$ neurons per hidden layer, and tanh activation. Unless otherwise specified, we use the BFGS optimiser from Optim.jl [Optimjl2018] to train NNs, following the settings in [Badia2024, Badia2025adaptive]. When the number of DoFs is relatively small (e.g., fewer than 1,000), NNs can easily overfit, so we reduce the number of training iterations (e.g., to a few hundreds). However, with sufficient number of DoFs, longer training does not negatively affect the training results, so we typically set a large number (e.g., 10,000) of iterations to ensure the NNs training process has converged. We also note that higher trial space orders require more iterations for the training to converge. Besides, by default, we use $\ell^{2}$-FEINNs (i.e., the Euclidean norm of the residual in loss function (7)), which has been proven to be effective for smooth solutions [Badia2024, Badia2025adaptive]. For rough solutions with localised high gradients, we use $H^{-1}$-FEINNs, i.e., loss (9) with $X=H^{1}$, which notably improves robustness as shown in Sect. 4.1.5.
To assess the accuracy of the identified solution $u^{id}$, we compute its $L^{2}$ and $H^{1}$ errors as:
$e_{L^{2}(\Omega)}(u^{id})=\left\lVert u-u^{id}\right\rVert_{L^{2}(\Omega)},\qquad e_{H^{1}(\Omega)}(u^{id})=\left\lVert u-u^{id}\right\rVert_{H^{1}(\Omega)},$
where $u$ is the true solution, $\left\lVert\cdot\right\rVert_{L^{2}(\Omega)}=\sqrt{\int_{\Omega}|\cdot|^{2}}$, and $\left\lVert\cdot\right\rVert_{H^{1}(\Omega)}=\sqrt{\int_{\Omega}|\cdot|^{2}+|\boldsymbol{\nabla}(\cdot)|^{2}}$. After training, where the interpolated NN (i.e., $\pi_{h}(u_{\mathcal{N}})$) is used to compute the loss and gradients, we obtain a NN solution (i.e., $u_{\mathcal{N}}$), $u^{id}$ can be either $\pi_{h}(u_{\mathcal{N}})$ or $u_{\mathcal{N}}$. For the $h$- and $p$-refinement error convergence tests, we repeat each experiment 10 times with different initial NN parameters to account for the sensitivity of achieved results with respect to NN initialisation. Since the variations in the $\pi_{h}(u_{\mathcal{N}})$ solutions are negligible, we only show the average errors among these 10 runs in the plots.
To establish a baseline to compare the accuracy of the identified solutions, we interpolate the true solution onto the trial FE space and compute the $L^{2}$ and $H^{1}$ errors. These interpolations are denoted as $\pi_{h}(u)$ in the figures and their errors are referred to as $\left\lVert u-\pi_{h}(u)\right\rVert$ in the text. There are several reasons that prevent us from using FEM solutions as a baseline for comparison. Firstly, the FEM solutions are sensitive to the choice of Nitsche coefficient $\gamma$. Secondly, unfitted FEMs often require additional stabilisation techniques, such as ghost penalties in CutFEM, which introduce extra parameters that can affect solution accuracy. Lastly, as mentioned in Sect. 3.1, the Petrov-Galerkin FE formulation proposed in this work leads to a rectangular system that cannot be directly solved using standard sparse direct solvers. In contrast, computing the FE interpolation of the true solution is straightforward and unaffected by these hyperparameters.
For the inverse experiments, we also utilise the BFGS optimiser. As recommended in [Badia2024], lighter NN structures are adopted, the activation function is set to softplus, and the $\ell^{1}$ norm of the residual vector is employed in the second term of the loss function (10). We follow the three-step training strategy proposed in [Badia2024, Sect. 4] to train the NNs. In step 1, the state NN $u_{\mathcal{N}}$ is trained using only the data misfit error (first term in (10)). In step 2, the unknown model parameter NNs $\boldsymbol{\Lambda}_{\mathcal{N}}$ are trained using the PDE residual (second term in (10)) with $u_{\mathcal{N}}$ fixed. Finally, in step 3, both $u_{\mathcal{N}}$ and $\boldsymbol{\Lambda}_{\mathcal{N}}$ are jointly trained using the full loss (10). The first two steps serve as cost-efficient initialisations for the final step, which is the most computationally intensive. Besides, step 3 is divided into several substeps, with the penalty coefficient $\alpha$ gradually increased to ensure the convergence of the training. For more details on the training strategy, refer to [Badia2024] and [Nocedal2006, Chap. 17].
4.1. 2D forward problems
As shown in Fig. 2, we consider two 2D geometries: a disk and a flower. The disk’s level-set function is $\phi(x,y)=\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}-R$, where $(x_{0},y_{0})$ is the centre and $R$ is the radius. The flower geometry, adapted from [Neiva2021], is described by $\phi(\rho,\theta)=\rho-0.12(\sin(5\theta)+3)$, where $\rho=\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}$ and $\theta$ is the polar angle relative to the centre $(x_{0},y_{0})$. We use the flower shape in Sect. 4.1.6 and the disk shape in the remaining subsections.
((a)) disk
((b)) flower
Figure 2. Analytic geometries used in the 2D experiments.
4.1.1. The effect of Nitsche coefficient for Poisson equation
In FEM, the choice of Nitsche coefficient $\gamma$ significantly affects solution accuracy. Thus, we first investigate how different $\gamma$ values impact the training of $\ell^{2}$-FEINNs for solving the Poisson equation. The physical domain is a disk centred at $(0.5,0.5)$ with a radius of $R=0.4$, and the background mesh is a $50\times 50$ uniform grid over $[0,1]^{2}$. The trial FE space has an order of $k_{U}=3$. We use different $\gamma$ values in the loss function and train NNs starting from the same initial parameters for each $\gamma$. We adopt the following smooth solution from [Badia2024]:
$u(x,y)=\sin(3.2x(x-y))\cos(x+4.3y)+\sin(4.6(x+2y))\cos(2.6(y-2x)).$ (11)
Fig. 3 shows the $L^{2}$ and $H^{1}$ error histories for the interpolated NNs and the NNs themselves during training with different values of $\gamma$ in the loss function. An immediate observation is that a large $\gamma$ value (e.g., $\gamma=10^{2}$) leads to a much slower convergence. However, convergence accelerates when $\gamma$ is small, and the corresponding $\pi_{h}(u_{\mathcal{N}})$ solutions eventually converge to the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. Another important finding is that, consistent with observations in [Badia2024, Badia2025adaptive], the NNs themselves can outperform their FE interpolations in terms of $L^{2}$ and $H^{1}$ errors. As shown in the second column of Fig. 3, the curves for $\pi_{h}(u_{\mathcal{N}})$ converge to the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, while the $u_{\mathcal{N}}$ curves eventually fall below these lines when $\gamma$ is suitable.
Figure 3. Convergence history of $u^{id}$ in $L^{2}$ and $H^{1}$ errors for the 2D Poisson problem, using different Nitsche coefficients during training. The top row shows $L^{2}$ errors and the bottom row represents $H^{1}$ errors. The first column displays errors of the interpolated NNs and the second column corresponds to errors of the NNs themselves. The red dashed lines represent the FEM errors.
Among the tested values, $\gamma=10^{-2}$ achieves the fastest convergence, so for convenience, we simply select $\gamma=10^{-2}$ for the remaining experiments in this work. However, we emphasise that $\ell^{2}$-FEINNs are robust to variations in $\gamma$ across a wide range. As shown in Fig. 3, the differences among convergence curves for $\gamma\leq 10$ values are not significant. For instance, $\gamma=10^{-4}$ and $\gamma=10$ yield similar error histories. Therefore, as long as $\gamma$ is “small enough” we expect $\ell^{2}$-FEINNs to perform well.
4.1.2. Comparison of PINNs and unfitted FEINNs
This experiment compares the computational efficiency of unfitted FEINNs with PINNs. The loss function for PINNs [Raissi2019] applied to the Poisson equation (1) reads:
$\mathscr{L}(u_{\mathcal{N}})=\frac{1}{N_{\rm int}}\sum_{i=1}^{N_{\rm int}}\left[\boldsymbol{\Delta}u_{\mathcal{N}}(x_{i},y_{i})+f(x_{i},y_{i})\right]^{2}+\tau_{b}\frac{1}{N_{\rm b}}\sum_{i=1}^{N_{\rm b}}\left[g(x_{i},y_{i})-u_{\mathcal{N}}(x_{i},y_{i})\right]^{2},$
where $N_{\rm int}$ and $N_{\rm b}$ are the number of interior and boundary collocation points, respectively, and $\tau_{b}$ is a penalty coefficient. The first term in the loss imposes the PDE, while the second term weakly enforces the Dirichlet boundary condition through penalisation.
We adopt the same disk geometry and smooth analytic solution as in Sect. 4.1.1. For the $\ell^{2}$-FEINN experiments, the background mesh consists of $25\times 25$ quadrilaterals, with $k_{U}=4$ and $\gamma=10^{-2}$. To ensure a fair comparison, we adopt the strategy in [Berrone2022], where PINNs use the same number of collocation points as the number of interpolation nodes employed by FEINNs. In particular, we extract the FEINN interpolation nodes and use those inside $\Omega$ as interior collocation points for PINNs. Using the level-set function, we then uniformly generate boundary collocation points such that their total number equals to the number the interpolation nodes located outside $\Omega$ in FEINNs. In this setup, $N_{\rm int}=5013$ and $N_{\rm b}=868$. This approach ensures both methods are trained on the same interior points and the same amount of total points. However, the computational cost is more favourable to PINNs, because as opposed to Nitsche’s method in FEINNs, they do not require computing spatial derivatives on the boundary (approximately 15% of the total points).
As suggested in [Berrone2022], convergence in PINN training can be accelerated by initially optimising with Adam and then switching to BFGS. However, convergence curves in [Berrone2022, Fig. 5] show that BFGS yields a steeper descent than Adam. Based on this observation and our preliminary experimental results, we select only BFGS for PINNs training. For the penalty coefficient $\tau_{b}$, we tested several values ranging from $10^{-4}$ to $10^{4}$ and found that $\tau_{b}=10^{3}$ yielded the best results for this problem. We implement PINNs in the TensorFlow framework [tensorflow2015]. For a fair comparison, $\ell^{2}$-FEINNs are also implemented and trained in TensorFlow.111 We are aware that PINNs can achieve higher accuracy with alternative optimisers, such as self-scaled BFGS and Broyden methods proposed in [Urban2025]. We also observe that different implementations of the same optimiser can yield different convergence behaviours. For instance, while SciPy’s BFGS can yield higher accuracy for PINNs with smaller NN architectures, it becomes too inefficient to use for larger NNs as the ones used in this work. A detailed exploration of these optimisers and the subtle differences in their implementations is beyond the scope of this work. All experiments are conducted on a single NVIDIA V100 GPU with 32 GB of memory. The $H^{1}$ error convergence history of NNs trained with PINNs and $\ell^{2}$-FEINNs is shown in Fig. 4.
Figure 4. Convergence history of the NN $H^{1}$ error over training iterations and wall-clock time in seconds for the 2D Poisson problem using PINNs and $\ell^{2}$-FEINNs. Both methods are implemented in TensorFlow, employ identical NN architectures with the same initialisation, and are trained on GPU using TensorFlow’s BFGS with an equal number of training points.
Let us first focus on the left subfigure of Fig. 4, which shows the convergence history of the NN $H^{1}$ error over BFGS iterations. When the penalty coefficient $\tau_{b}$ is appropriately chosen (e.g., $\tau_{b}=10^{3}$), PINNs achieve accuracy comparable to $\ell^{2}$-FEINNs in terms of $H^{1}$ error. However, the choice of $\tau_{b}$ significantly affects PINNs’ accuracy: the curve for $\tau_{b}=10^{-1}$ is placed well above that for $\tau_{b}=10^{3}$, indicating that an unsuitable $\tau_{b}$ can severely hinder PINN convergence. In contrast, as shown in Fig. 3, NNs trained by $\ell^{2}$-FEINNs converge below the $\left\lVert u-\pi_{h}(u)\right\rVert$ reference lines even with suboptimal $\gamma$ values, demonstrating unfitted FEINNs’ greater robustness to hyperparameter variation.
The right subfigure of Fig. 4 compares the computational time for training PINNs and $\ell^{2}$-FEINNs. Notably, the $\ell^{2}$-FEINN method completes 5,000 BFGS iterations in around 750 seconds, while PINNs requires around 1,400 seconds. This substantial efficiency gain of $\ell^{2}$-FEINNs stems not only from avoiding computation of spatial derivatives of NNs to accelerate loss and gradient evaluations, but also from a better-posed optimisation problem that allows BFGS to converge with fewer gradient calculations. Specifically, $\ell^{2}$-FEINNs require 11,308 gradient computations, compared to 13,031 for PINNs with $\tau_{b}=10^{3}$. Moreover, each gradient computation takes 0.0645 seconds for $\ell^{2}$-FEINNs, versus 0.1064 seconds for PINNs. These results align with the findings in [Berrone2022], where IVPINNs were also shown to be faster than PINNs and VPINNs in training time.
In summary, for smooth solutions, PINNs have the potential to achieve similar accuracy as unfitted $\ell^{2}$-FEINNs; however, their performance is sensitive to the choice of the hyperparameter $\tau_{b}$. Besides, by avoiding the computation of spatial and nested derivatives of NNs, $\ell^{2}$-FEINNs are faster. It is important to mention that, in this experiment, the level-set function enables easy generation of collocation points for PINNs. However, for more complex geometries, such as those defined by STL meshes, as discussed in Sect. 4.2.2, generating collocation points becomes challenging. In contrast, $\ell^{2}$-FEINNs can handle such complex geometries without additional effort, as demonstrated in the 3D experiments.
4.1.3. Moving domain test for the Poisson equation
We now examine the robustness of unfitted $\ell^{2}$-FEINNs to changes in the relative position between the physical domain and the background mesh. We adapt the moving disk experiment from [Badia2018, Sect. 6.2], where a disk with radius $R=0.19$ moves along the anti-diagonal of the background box $[0,1]^{2}$ by shifting its centre $(x,x)$. As $x$ varies, the relative position of the disk with respect to the background mesh changes, leading to a comprehensive exploration of different cut cell configurations. The goal of the experiment is to study how the accuracy of the trained NNs might be affected by small cut cells. We use $k_{U}=3$ and a uniform background mesh with $50\times 50$ quadrilaterals. The value of $x$ ranges from 0.2 to 0.8, with a small step size of 0.005 to increase the probability of encountering small cut cells. The true solution is still given by (11).
Fig. 5 displays the $L^{2}$ and $H^{1}$ errors of the identified solution versus the centre position of the moving disk. The NNs start with the same initial parameters for each $x$ value. For the interpolated NNs, both $L^{2}$ and $H^{1}$ errors closely follow the flat $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. While there are fluctuations in the NN errors, they consistently remain below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. Notably, the $H^{1}$ errors of the NNs are steadily more than an order of magnitude lower than the $\pi_{h}(u_{\mathcal{N}})$ errors. These observations highlight that the unfitted $\ell^{2}$-FEINN method is highly robust to small cut cells without using any kind of FE stabilisation such as AgFEM or CutFEM to compute the residual. In the context of unfitted FEINNs, such stabilisations are required only when using the discrete dual norm in the loss function (9), where the discrete dual norms of the residual must be computed.
Figure 5. $L^{2}$ and $H^{1}$ errors of the trained NNs for the 2D Poisson problem plotted against the centre position of the moving disk.
4.1.4. Convergence tests for the Poisson equation
In previous experiments, we have established that a Nitsche coefficient of $\gamma=10^{-2}$ yields accurate solutions for $k_{U}\in\{2,3\}$ and confirmed the robustness of unfitted $\ell^{2}$-FEINNs against small cut cells. We now confidently proceed to investigate the convergence properties of $\ell^{2}$-FEINNs with respect to the trial space background mesh size ($h$) and order ($k_{U}$). We use the disk geometry centred at $(0.5,0.5)$ with radius $R=0.4$ and a uniform Cartesian background mesh on $[0,1]^{2}$. The true solution is again (11).
We first fix $k_{U}$ and perform an $h$-refinement experiment for unfitted $\ell^{2}$-FEINNs. The convergence results are shown in Fig. 6. Different colours distinguish between $k_{U}=2$ and $k_{U}=3$, while markers are used to differentiate between NNs and their interpolations. The dashed lines represent errors of $\pi_{h}(u)$ versus mesh size.
Figure 6. Convergence of the $L^{2}$ and $H^{1}$ errors with respect to the trial space mesh size for the 2D Poisson problem. The figure shows results for two orders: $k_{U}=2$ and $k_{U}=3$, distinguished by colours.
All the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines in Fig. 6 exhibit the expected slopes, which validates the accuracy of the analytic solution FE interpolation and confirms that the error computations use sufficient Gaussian quadrature degree. The $\pi_{h}(u_{\mathcal{N}})$ errors always align with the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines for both $k_{U}=2$ and $k_{U}=3$, indicating that $\ell^{2}$-FEINNs converge at the optimal rate. Both $L^{2}$ and $H^{1}$ errors for NNs are below the corresponding $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, manifesting NNs’ superior accuracy compared to their interpolation counterparts and $\pi_{h}(u)$. Remarkably, the finest mesh for $k_{U}=3$ yields $H^{1}$ errors that are about two orders of magnitude lower than the $\pi_{h}(u_{\mathcal{N}})$ errors. Additionally, the impact of NN initialisation is minimal, as the NN errors from different initialisations are closely clustered.
Given the smoothness of the analytic solution (11), we also investigate how $k_{U}$ affects the convergence rate of unfitted $\ell^{2}$-FEINNs through $p$-refinement tests. We fix the background mesh size and increase $k_{U}$ from 1 to 5. The experiment results are displayed in Fig. 7. We consider two mesh sizes, $h=1/14$ and $h=1/21$, differentiated by colours in the figure. Similar to the $h$-refinement test, the $\pi_{h}(u_{\mathcal{N}})$ errors agree with the true solution interpolation errors, indicating that the method accomplishes the expected convergence rate as $k_{U}$ increases for both mesh sizes. The NN errors are positioned below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, demonstrating the improved NN accuracy without interpolations. Additionally, Fig. 7 validates the choice of $\gamma=10^{-2}$ for $k_{U}$ values ranging from 1 to 5, building on the previous subsections where $k_{U}\in\{2,3\}$.
Figure 7. Convergence of the $L^{2}$ and $H^{1}$ errors with respect to the trial space order for the 2D Poisson problem. The figure shows results for two mesh sizes: $h=1/14$ and $h=1/21$, distinguished by colours.
In [Badia2024, Sect. 6.1.1], we also conducted $h$- and $p$-refinement tests for the body-fitted FEINN method on the convection-diffusion-reaction equation, where Dirichlet boundary conditions were strongly imposed via FE offset functions. The results showed that FEINNs achieved the expected convergence rates in both tests, and NNs outperformed FEM approximations when the analytic solution was smooth. The current experiments yield similar results for the Poisson equation, further confirming the reliability and accuracy of the FEINN framework for non-trivial geometries, even when Dirichlet boundary conditions are weakly imposed through Nitsche’s method.
4.1.5. Discrete dual norms for the Poisson equation
In this experiment, we explore the impact of discrete dual norms in the training of unfitted FEINNs for the Poisson equation. We compare the performance of PINNs with both unfitted $\ell^{2}$- and $H^{-1}$-FEINNs. We use the same disk geometry as in Sect. 4.1.4 and set $k_{U}=3$. The background box $[0,1]^{2}$ is partitioned into $50\times 50$ quadrilaterals. To increase the difficulty of the problem, we use the following analytic solution, which features a sharp gradient around the centre of the disk:
$u(x,y)=\sin(3.2x(x-y))(5\mathrm{e}^{-100((x-0.5)^{2}+(y-0.5)^{2})}+1).$
Since the solution is rougher than (11), the use of a discrete dual norm is instrumental in improving the convergence of the training process.
As described in Sect. 2.4, we consider two linearised test FE spaces: the standard active space $V_{h}^{\rm{std}}$ consisting of all internal and cut elements, and the aggregated active space $V_{h}^{\rm{ag}}$ The discrete dual norm requires to compute the Gram matrix for the linearised test space with respect to the $H^{1}$-inner product. However, spectrally equivalent choices can also be considered. For example, in [Badia2024, Sect. 6.1.3], we demonstrated that a single cycle of GMG was an effective approximation of the inverse of the Gram matrix. Since small cut cells in $V_{h}^{\rm{std}}$ may cause the Gram matrix to become singular in the presence of rounding errors, additional ghost penalty terms (5) are required to ensure the Gram matrix is well-posed. Note that in the context of unfitted FEINNs, even though the trial space may be of higher order, $V_{h}^{\rm{std}}$ is always linear. Therefore, a first-order penalty term is sufficient to stabilise the Gram matrix, i.e., $k_{U}=1$ in (5). The implementation of unfitted FEINNs does not actually need to deal with high order derivatives in the case of CutFEM stabilisation, nor specialised discrete extensions required for high-order aggregated FE spaces [Badia2022c]. Following the recommendation in [Burman2015], we set $\gamma_{g}=0.1$ in this experiment. On the other hand, the $V_{h}^{\rm{ag}}$ space is designed to overcome stability issues by construction, thus no additional penalty terms are needed in the weak formulation.
As shown earlier in Fig. 3, large Nitsche coefficients can slow down the training convergence, so we include both the suitable $\gamma=10^{-2}$ and a large $\gamma=10^{2}$ in the experiment to test how effectively $H^{-1}$-FEINNs handles unsuitable Nitsche coefficients. For clarity in the plots, in this experiment, we only consider the $H^{1}$ primal norm in the loss (9). However, other primal norms can also be considered. For example, in [Badia2025adaptive, Sect. 5.1.2], we showed that four different primal norms were effective for the Poisson equation.
We plot the error histories for $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ for different combinations of test spaces for $\ell^{2}$-FEINNs and $H^{-1}$-FEINNs in Fig. 8. In the figure, we label $\ell^{2}$-FEINNs with the postfix “$V_{h}^{\rm{std}}$, $\ell^{2}$” and $H^{-1}$-FEINNs with the postfixes “$V_{h}^{\rm{std}}$, $H^{-1}$” or “$V_{h}^{\rm{ag}}$, $H^{-1}$”. We observe that for $\ell^{2}$-FEINNs, a large Nitsche coefficient also impedes the convergence of both $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ for this problem (with a solution featuring a sharp gradient around the disk centre). However, the use of dual discrete norms in $H^{-1}$-FEINNs overcomes this issue and accelerates the convergence significantly. For instance, with $\gamma=10^{2}$, the $L^{2}$ curve of $\pi_{h}(u_{\mathcal{N}})$ for $H^{-1}$-FEINNs drops to $10^{-5}$ at around 800 iterations, while $\ell^{2}$-FEINNs require more than 4,000 iterations to achieve the same error level. Besides, the use of a discrete dual norm in the loss function is also effective for the suitable Nitsche coefficient ($\gamma=10^{-2}$), as it does not not only accelerates the convergence of NNs but also enhances the NN solution accuracy. As shown in Fig. 8, the corresponding error curves decrease more rapidly than the $\ell^{2}$-FEINN ones. In addition, the NN error curves for $H^{-1}$-FEINNs in the second column of Fig. 8 drop below the $\ell^{2}$-FEINN ones shortly after training begins.
Figure 8. Convergence history of $u^{id}$ in $L^{2}$ and $H^{1}$ errors for the 2D Poisson problem, using different combinations of Nitsche coefficient, test space, and norm during training. Refer to the caption of Fig. 3 for details on the information being displayed in this figure.
It is clear from Fig. 8 that all the dual norm error curves for different $\gamma$ and test space combinations are very close. This implies that dual norm losses are robust to variations in $\gamma$ values and that the choice of test space has only a minor impact on their effectiveness. Since the Gram matrix requires neither additional stabilisation nor parameter tuning, we recommend using $V_{h}^{\rm{ag}}$ as the test space in practice. However, we reiterate that the unfitted FEINN framework only requires a first-order ghost penalty term to stabilise $V_{h}^{\rm{std}}$, which makes $V_{h}^{\rm{std}}$ also a practical choice.
In terms of computational cost, we note that the Gram matrix is symmetric positive definite and is built and factorised using sparse Cholesky factorisation only once before training. Since the problem is linear, this factorisation can be stored and reused throughout training. At each iteration, the inverse of the Gram matrix is applied to the residual vector by forward and backward substitutions using the sparse Cholesky factor, which makes the training computationally efficient.
[…]
In Fig. 9, we present NN error history versus training iterations and time in seconds. Let us first comment on the convergence curves of $\ell^{2}$-FEINNs and $H^{-1}$-FEINNs. The use of dual norms indeed improves computational efficiency. In the $L^{2}$ error plot on the left of Fig. 9, even with a suitable Nitsche coefficient ($\gamma=10^{-2}$), $\ell^{2}$-FEINNs require substantially more training time than $H^{-1}$-FEINNs to reach comparable accuracy. For instance, achieving the reference $\left\lVert u-\pi_{h}(u)\right\rVert$ error, as indicated by the red dashed lines, takes around 2,000 iterations and 3,000 seconds for $\ell^{2}$-FEINNs, compared to just 800 iterations and 1,000 seconds for $H^{-1}$-FEINNs. Another notable observation is that the overall computational cost of $\ell^{2}$- and $H^{-1}$-FEINNs is similar, with both methods finishing 3,000 iterations in around 4,000 seconds. This suggests that incorporating dual norms does not significantly increase the computational cost, while still accelerating convergence.
In the PINN experiments, we again tested different $\tau_{b}$ values and found that $\tau_{b}=10^{4}$ generally yielded the best results. Among the three methods, PINNs exhibit the slowest convergence in both time and iterations, requiring more than 2,700 iterations and 7,800 seconds to reach the reference $\left\lVert u-\pi_{h}(u)\right\rVert$ error; $H^{-1}$-FEINNs is noticeably the most robust method.
[…]
4.1.6. Convergence test for a nonlinear equation
As a final forward 2D experiment, we solve the nonlinear equation (2) on the flower geometry (see Fig. 2(b)), with coefficients $\boldsymbol{\beta}=[2,3]^{\rm T}$ and $\sigma=4$. The flower is centred at $(0.5,0.5)$ within the background box $[0,1]^{2}$. The true solution is
$u(x,y)=\frac{\cos(\pi(3x+y))}{1+(x+2y)^{2}}.$
We use the $\ell^{1}$ norm in the loss (7) for this experiment.
Similar to Sect. 4.1.4, we conduct $h$-refinement tests for the nonlinear equation on the flower domain. We use $k_{U}\in\{2,3\}$ and change the background mesh size for each $k_{U}$. The results are shown in Fig. 10. For this more challenging nonlinear problem, the $\pi_{h}(u_{\mathcal{N}})$ errors still exhibit the expected convergence rates, aligning with the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. The NN errors constantly fall below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, suggesting a similar accuracy improvement over $\pi_{h}(u_{\mathcal{N}})$ as seen in the Poisson equation experiments.
Figure 10. Convergence of the $L^{2}$ and $H^{1}$ errors with respect to the trial space mesh size for the 2D nonlinear problem. The figure shows results for two orders: $k_{U}=2$ and $k_{U}=3$, distinguished by colours.
4.2. 3D forward problems
We research two analytic geometries in 3D: a doughnut and a popcorn, as illustrated in Fig. 11. Their level-set surface expressions are detailed in [Burman2015]. For the doughnut, we choose $R=0.3$ and $r=0.1$; for the popcorn, we select $r_{0}=0.3$, $\sigma=0.09$, and $A=0.9$. Both geometries are centred at $(0.5,0.5,0.5)$. To further demonstrate the capability of unfitted FEINNs in handling complex 3D geometries, we also include an octopus geometry described explicitly using an STL mesh. We apply the $\ell^{1}$ norm in the loss (7) for all 3D experiments.
((a)) doughnut
((b)) popcorn
Figure 11. Analytic geometries used in the 3D experiments.
4.2.1. Convergence tests on analytic geometries
We first conduct $h$-convergence tests on the doughnut and popcorn geometries. We use trial FE space order of $k_{U}=2$ and alter the mesh resolution for the background box $[0,1]^{3}$. The true solution for the Poisson equation is
$u(x,y,z)=\sin(4\pi\sqrt{(x-1)^{2}+(y-1)^{2}+(z-1)^{2}}).$
Fig. 12 illustrates the decay of $L^{2}$ and $H^{1}$ errors for $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ as the mesh size changes, with different colours representing the two geometries. For reference, the $\left\lVert u-\pi_{h}(u)\right\rVert$ dashed lines, showing the expected slopes, are also included. The $\pi_{h}(u_{\mathcal{N}})$ errors clearly achieve the expected convergence rates as they all fall on the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. The NN errors are well below the dashed lines, suggesting superior accuracy of the NNs, even with complex 3D geometries defined by level-set functions.
Figure 12. Convergence of the $L^{2}$ and $H^{1}$ errors with respect to the trial space ($k_{U}=2$) mesh size for the 3D Poisson problem. The figure shows results for two geometries: a popcorn and a doughnut, distinguished by colours.
4.2.2. Training on the octopus geometry defined by STL meshes
As a concluding experiment for forward problems, we focus on a typical computer-aided engineering (CAE) scenario where the boundary of complex geometries are described by surface STL meshes. In practice, the analytic level-set function of a geometry is often unavailable, and the STL mesh is a reliable alternative for providing geometric information. We adopt the octopus geometry, which consists of 10,173 vertices and 20,342 faces, from the Thingi10K dataset [Zhou2016thingi10k]. This geometry, identified as file ID 69930 in the dataset, is depicted in Fig. 13. We utilise the Gridap.jl ecosystem’s Julia package, STLCutters.jl [Martorell2024], that provides algorithms to compute the intersection of STL meshes and background meshes [Badia2022b]. The trial FE space of order $k_{U}=3$ is constructed on a $40\times 40\times 20$ quadrilateral background mesh over the box $[-80,80]\times[-80,80]\times[-10,40]$. We solve the Poisson problem, with the true solution illustrated in Fig. 13(a) and defined by
$u(x,y,z)=\cos(\sqrt[3]{(x-100)^{2}+(y-100)^{2}+(z-50)^{2}}).$
((a)) $u$
((b)) $|u-u_{\mathcal{N}}|$
((c)) $|\boldsymbol{\nabla}(u-u_{\mathcal{N}})|$
Figure 13. Reference solution and NN errors for the 3D Poisson problem on the octopus geometry.
We present the NN error in Fig. 13(b) and its gradient error magnitude in Fig. 13(c). The NN solution is highly accurate, with point-wise errors below $3\times 10^{-5}$ and gradient error magnitudes below $6\times 10^{-5}$. For comparison, Tab. 1 lists the errors for the FE interpolation of the analytic solution $u$, $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ solutions. As expected, the $\pi_{h}(u_{\mathcal{N}})$ errors are comparable to the $\pi_{h}(u)$ errors, while the NN errors are notably smaller than those of $\left\lVert u-\pi_{h}(u)\right\rVert$. Tab. 1 confirms that unfitted FEINNs are able to accurately approximate the Poisson problem on complex 3D geometries. It also demonstrates that, for smooth solutions, the trained NN can achieve higher accuracy than its interpolation, even with highly intricate 3D shapes defined by STL meshes. Additionally, it also opens the possibility of using FEINNs for solving problems in real-world CAE applications.
$\pi_{h}(u)$ $\pi_{h}(u_{\mathcal{N}})$ $u_{\mathcal{N}}$
$L^{2}$ error $1.78\times 10^{-3}$ $1.51\times 10^{-3}$ $6.33\times 10^{-4}$
$H^{1}$ error $4.53\times 10^{-3}$ $4.41\times 10^{-3}$ $6.94\times 10^{-4}$
Table 1. Errors of the FE interpolation of the true solution, interpolation of the trained NN and NN itself for the 3D Poisson problem on the octopus geometry.
4.3. Inverse nonlinear problems
We conclude the numerical experiments by solving a nonlinear inverse problem on the flower geometry in Fig. 2(b). The objectives are to estimate the coefficient $\sigma$ in the nonlinear equation (2) from partial observations of the state $u$, and to fully reconstruct the partially known $u$. We set $\boldsymbol{\beta}=[2,3]^{\mathrm{T}}$ and define appropriate $f$ and $g$ such that the analytical state and coefficient are given by
$u(x,y)=\sin(\pi x)\sin(\pi y),\quad\sigma(x,y)=1+9\mathrm{e}^{-5((x-0.5)^{2}+(2y-1)^{2})}.$
The first column of Fig. 14 displays the true state and coefficient. The observed data are limited to the central region of the domain, indicated by purple dots in Fig. 14(a). Note that we assume prior knowledge of the sign of the unknown coefficient $\sigma$, i.e., $\sigma>0$.
((a)) $u$
((b)) $u_{\mathcal{N}}$
((c)) $|u-u_{\mathcal{N}}|$
((d)) $\sigma$
((e)) $\sigma_{\mathcal{N}}$
((f)) $|\sigma-\sigma_{\mathcal{N}}|$
Figure 14. Illustration of analytical (left), NN solutions (center) and errors (right) for the inverse nonlinear problem with partial observations. The purple dots in (a) represent the observed data.
We adopt the NN architectures from [Badia2024, Sect. 6.2.1]: both NNs $u_{\mathcal{N}}$ and $\sigma_{\mathcal{N}}$ have two hidden layers with 20 neurons each; the output layer of $\sigma_{\mathcal{N}}$ is equipped with a rectification function $r(x)=|x|+0.01$ to ensure positivity. All the FE spaces, i.e., $u_{\mathcal{N}}$ interpolation or trial space, $\sigma_{\mathcal{N}}$ interpolation space, and test space, are continuous Galerkin spaces of first order on a background mesh consisting of $50\times 50$ quadrilaterals. These FE spaces are standard unfitted spaces without any stabilisation. Similar to forward problems, essential boundary conditions are weakly imposed via Nitsche’s method with $\gamma=10^{-2}$. The training starts with 400 iterations for $u_{\mathcal{N}}$ in step 1, followed by 100 iterations for $\sigma_{\mathcal{N}}$ in step 2. Step 3 compromises three substeps, with 500 iterations each for the coupled training of $u_{\mathcal{N}}$ and $\sigma_{\mathcal{N}}$. The penalty coefficients $\alpha$ are set to 0.01, 0.03 and 0.09 for these substeps, respectively.
The last two columns of Fig. 14 show the NN solutions and their point-wise errors. Even though the state observations are limited to partial regions of the domain, the trained $u_{\mathcal{N}}$ in Fig. 14(b) accurately approximates the true state in Fig. 14(a) across the entire domain, with point-wise errors below $6\times 10^{-4}$ as shown in Fig. 14(c). The trained $\sigma_{\mathcal{N}}$ in Fig. 14(e) also accurately captures the the true coefficient in Fig. 14(d) in most regions of the domain, with point-wise errors in Fig. 14(f) mostly concentrated around certain parts of the domain boundary. The errors are below $0.2$ in most regions, with the maximum reaching $0.76$ near the boundary.
The relative errors during training for the state and coefficient solutions corresponding to the results in Fig. 14 are displayed in Fig. 15. The label “NN interp” refers to the interpolated NN solutions, while “NN” refers to the NN solutions themselves. The definition of the relative error is the norm of the difference between the true and identified solutions, divided by the norm of the true solution. It is evident from Fig. 15 that the three-step training strategy effectively addresses the inverse nonlinear problem with partial observations. The state solution errors decrease rapidly during step 1 (iterations 1-400), while the coefficient solution errors decrease quickly during step 2 (iterations 401-500). The coupled training in step 3 further reduces the errors in both state and coefficient solutions. For the state, the interpolated NN solution $\pi_{h}(u_{\mathcal{N}})$ stops improving after 700 iterations due to the low-order interpolation of $u_{\mathcal{N}}$. Despite this, $\pi_{h}(u_{\mathcal{N}})$ already provides satisfactory solution, with relative $L^{2}$ and $H^{1}$ errors below $0.1\%$ and $3\%$, respectively. Moreover, the state NN solution $u_{\mathcal{N}}$ continues to improve, achieving an order of magnitude gain in accuracy in terms of the relative $H^{1}$ error. For the coefficient, both the NN solution $\sigma_{\mathcal{N}}$ and its interpolation $\pi_{h}(\sigma_{\mathcal{N}})$ exhibit similar convergence curves, with their relative $L^{2}$ errors steadily decreasing to $2\%$.
Figure 15. Relative errors during training by the unfitted FEINN method for the inverse nonlinear problem with partial observations.
To further evaluate the robustness of unfitted FEINNs against varying NN initialisations, we repeat the experiment 100 times using different initial parameters for both $u_{\mathcal{N}}$ and $\sigma_{\mathcal{N}}$. The box plots of the resulting relative errors are displayed in Fig. 16. The interpolated NN solutions for the state are consistently accurate across various NN initialisations, with almost flat box plots centred at 0.07% and 3% for the relative $L^{2}$ and $H^{1}$ errors, respectively. Although, the NN solutions for the state show larger variation compared to the interpolations, their boxes are still relatively compact. Both the relative $L^{2}$ and $H^{1}$ error boxes are positioned below those of the interpolated NN solutions, indicating improved accuracy. Notably, for the relative $H^{1}$ error, the third quartile falls below 0.3%, an order of magnitude lower than the corresponding interpolated NN solutions. The NN solutions and their interpolations for the coefficient are nearly identical, with relative $L^{2}$ error showing first and third quartiles at 2% and 4%, respectively. Overall, Fig. 16 confirms the robustness of unfitted FEINNs with respect to different NN initialisations in inverse problems with partial observations. This experiment further extends the FEINN framework from linear inverse problems [Badia2024, Sect. 6.2] to nonlinear ones, showcasing the method’s high accuracy and reliability on complex geometries.
Figure 16. Box plots from 100 experiments for the inverse nonlinear problem with partial observations.
5. Conclusions
We combine unfitted FE discretisations and the FEINN framework for the approximation of PDEs on complex geometries using NNs. In the proposed unfitted FEINN method, Dirichlet boundary conditions are weakly imposed via Nitsche’s method. The high-order trial FE space onto which the NN is interpolated is built on the active cells of a coarse background mesh that is unfitted to the boundary of the geometry. With ease of implementation, and existing unfitted FEM code reuse in mind, the linearised test space is constructed on the active cells of a uniformly refined version of this coarse background mesh. Although the trial space has more DoFs than the test space, which leads to a rectangular system, this mismatch can be naturally accommodated by the FEINN method as such framework is based on a residual minimisation principle. Indeed, unfitted FEINNs exhibit outstanding convergence and accuracy in practice.
We conduct extensive numerical experiments to evaluate the unfitted FEINN method on various 2D and 3D complex geometries for different values of numerical parameters, geometries and PDEs. As opposed to unfitted FEM, small cut cells have negligible impact on the accuracy of both trained NNs and their interpolations, even without using stabilisation techniques such as CutFEM or AgFEM to compute the residual (such stabilisations are required, though, for the computation of discrete dual norms). The $h$- and $p$-convergence tests for the Poisson equation manifest that the trained NNs outperform the FE interpolation of the analytic solution on the same trial space, and their interpolations accomplish the expected convergence rates. Besides, unfitted FEINNs demonstrate robustness to the choice of Nitsche coefficient, which is a another advantage over unfitted FEMs.
In experiments comparing unfitted FEINNs and PINNs, FEINNs achieve comparable or better accuracy, demonstrate higher robustness to hyperparameter choices, and require significantly less training time on both GPUs and CPUs. We observe that the use of a discrete dual norm of the residual in the loss function, combined with a robust unfitted discretisation, significantly accelerates NN training convergence compared to the Euclidean norm variant and PINNs. This effect is particularly notable for Poisson problems featuring a sharp gradient in a localised region of the domain, substantially reducing computational cost. The nonlinear equation experiments verify that unfitted FEINNs are capable of dealing with more challenging PDEs. The experiments on 3D geometries show that unfitted FEINNs accurately solve PDEs on 3D complex geometries defined implicitly via level-set functions or explicitly through STL surface meshes. This makes the method well-suited for real-world CAE scenarios.
The inverse problem experiments exhibit that unfitted FEINNs can be easily and effectively extended to solve nonlinear inverse problems, delivering highly accurate and robust solutions on complex geometries. This overcomes key limitations of traditional FEMs for inverse settings, which require the use of adjoint solvers, nested loops, and regularisation.
Future research could explore several promising directions. First, as discussed in [Badia2025adaptive], to exploit the nonlinear approximability properties of NNs, it is essential to adaptively refine the FE spaces. Enhancing unfitted FEINNs with adaptive training techniques [Badia2025adaptive] will enable us to tackle problems with multiscale features or singularities on complex geometries. Another encouraging research avenue is applying unfitted FEINNs to evolutionary PDEs, especially time-dependent problems on moving domains [Badia2023] and fluid-structure interaction problems with evolving geometries [Massing2015]. In time-dependent problems, the NN trained at the previous time step serves as a good initial guess for the current step, which can clearly improve the training efficiency of unfitted FEINNs.
Acknowledgments
This research was partially funded by the Australian Government through the Australian Research Council (project numbers DP210103092 and DP220103160). This work was also supported by computational resources provided by the Australian Government through NCI under the NCMAS and ANU Merit Allocation Schemes. W. Li gratefully acknowledges the Monash Graduate Scholarship from Monash University, the NCI computing resources provided by Monash eResearch through Monash NCI scheme for HPC services, and the support from the Laboratory for Turbulence Research in Aerospace and Combustion (LTRAC) at Monash University through the use of their HPC Clusters.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
[…]